### v1

@@ I Introduction
Time series forecasting, especially in the financial sector, is a challenging, yet critically demanded and safety critical problem. Market prices are influenced by a multitude of factors, leading to complex investment and trading strategies that are required to model and manage multiple aspects [1]. In this paper, we focus on the particular problem of stock price prediction. The stock market can be viewed as a dynamic system, where stocks and further financial commodities (e.g. currencies, equities) are being traded. Traders generally prefer to purchase stocks of companies that show potential for future growth and avoid those expected to decline in value [2]. The most commonly used methods for predicting stock returns are broadly categorized into fundamental analysis and technical analysis. Fundamental analysis aims to determine the intrinsic value of stocks based on internal and external economic indicators, that is particularly essential for investors focused on long-term trading. Technical analysis, being investigated in this paper, focuses on techniques to maximize the capital gain and decrease the loss, based on statistical analysis that is demanded to forecast the stock price [3]. Fast and accurate stock price forecasting is difficult, and this field is an active area of research where the best ways are yet to be found [2].
@@ IV Results and Evaluation
Figure 5: The predicted close prices. NASDAQ (top) and S&P500 (bottom) predicted close prices and orginal close prices are displayed in the same interval of 6-month (from 2024-08-01 to 2025-02-05) together with their trends and predicted trends. The last date of data was used for training is 2024-09-30.
In Table II, we compared our model with different encoding techniques, and state-of-the-art deep learning predictor models, such as RNN and LSTM. Here, TT2VFin refers to our proposed multi-features model, TT2VFin+p adds positional encoding layer, TT2VFin+M uses masking in the attention mechanism, TT2VFin+p+M uses both positional encoding and masking. The results show that TT2VFin outperforms its variations. Applying a mask does not significantly affect TT2VFin performance; however, the positional encoding layer confuses the model and produces poor results. The proposed model outperforms RNN despite utilizing about three times fewer parameters as RNN. LSTM demonstrates its power in financial prediction; nonetheless, there is still a performance difference between TT2VFin and LSTM. Statistically, TT2VFin outperforms LSTM by around 9.25% $\pm$ 0.63% in Group A (averaging the metrics). We achieved the same results in other groups, where our proposed model is dominating over others. TT2VFin+p and TT2VFin+M have no change in performance, but TT2VFin+p+M cannot maintain its previous performance. In general, our proposed model uses 2 to 3 times fewer parameters than others but has roughly 14.23% greater performance (by averaging all the metrics differences between base and RNN and LSTM).
@@
In this experiment, TT2VFin and its variations use 143K parameters, RNN uses 559K parameters, and LSTM uses 355K parameters.
@@ V Conclusion
In this paper, we investigated the application of deep learning for stock price prediction, which is a challenging time series forecasting problem in the financial industry. We studied different markets by selecting correlation features to enhance the accuracy in predicting multiple stock prices and developed a novel neural network architecture that combines the Encoder of the Transformer with the Time2Vec model. Through careful testing and analysis, we have found that our model performs exceptionally well, better than standard models such as LSTM and RNN, and encoding methods like positional encoding from Transformer. In most cases, our multi-feature proposed model is better than a single-feature model, that verifies the importance and potential of feature selection and aggregation. In conclusion, our study demonstrated that techniques like Time2Vec and Encoder are promising future approaches combined with multi-source feature selection. We also believe that our approach can be successfully adopted and generalized to related financial and other time series prediction problems, and can achieve even better results for more refined data sources (e.g., having the intra-day historical data).

### v2

@@ I Introduction
Time series forecasting, especially in the financial sector, is a challenging, yet critically demanded and safety critical problem. Market prices are influenced by a multitude of factors, leading to complex investment and trading strategies that are required to model and manage multiple aspects [1]. In this paper, we focus on the particular problem of stock price prediction. The stock market can be viewed as a dynamic system, where stocks and further financial commodities (e.g. currencies, equities) are being traded. Traders generally prefer to purchase stocks of companies that show potential for future growth and avoid those expected to decline in value [2]. Companies within the same sector often exhibit similar correlations, as they tend to be affected similarly by macroeconomic events such as economic downturns or political policies. The most commonly used methods for predicting stock returns are broadly categorized into fundamental analysis and technical analysis. Fundamental analysis aims to determine the intrinsic value of stocks based on internal and external economic indicators, that is particularly essential for investors focused on long-term trading. Technical analysis, being investigated in this paper, focuses on techniques to maximize the capital gain and decrease the loss, based on statistical analysis that is demanded to forecast the stock price [3]. Fast and accurate stock price forecasting is difficult, and this field is an active area of research where the best ways are yet to be found [2].
@@ III-C Proposed Model
Our design concepts were the to develop a robust architecture that can utilize the expression and generalization ability of the attention mechanism through Encoders, and to involve Time2Vec as encoding. Our investigations proved that Time2Vec is a fitting encoding scheme for the target application, and provides better ability to capture temporal patterns compared to the traditional approaches, as discussed in Section IV.
@@ IV Results and Evaluation
In this experiment, TT2VFin and its variations use 143K parameters, Transformer+p uses 173K parameters, RNN uses 559K parameters, and LSTM uses 355K parameters.
Figure 5: The predicted close prices. S&P500 predicted close prices and orginal close prices are displayed in the same interval of 6-month (from 2024-08-01 to 2025-02-05) together with its trend and predicted trend. The red line demonstrates the last collected date we did in Section III-B. Note that a similar pattern of close prices and trends are attained by the NASDAQ stock.
In Table II, we compared our model with different encoding techniques, and state-of-the-art deep learning predictor models, such as RNN and LSTM. Here, TT2VFin refers to our proposed multi-features model, TT2VFin+p adds positional encoding layer, TT2VFin+M uses masking in the attention mechanism, TT2VFin+p+M uses both positional encoding and masking. The results show that TT2VFin outperforms its variations, and Transformer+p is a Transformer Encoder using positional encoding instead of Time2Vec. As a more detailed analysis, we can see that applying a mask does not significantly affect TT2VFin; however, the incorporation of positional encoding to the proposed model reduces its performance. The likely reason behind this behavior is that Time2Vec itself acts as an encoding scheme, and the combination of the two encodings confuses the model and leads to suboptimal performance. We also see that TT2VFin outperforms Transformer+p in 5 out of 6 cases and has comparable performance in one case, that proves that Time2Vec is more efficient for the time-series domain than the original positional encoding. Regarding the baseline deep learning models, the proposed model outperforms RNN despite utilizing about three times fewer parameters as RNN. LSTM demonstrates its power in financial prediction; nonetheless, there is still a performance difference between TT2VFin and LSTM. Statistically, TT2VFin outperforms LSTM by around 9.25% $\pm$ 0.63% in Group A (averaging the metrics). We achieved the same results in other groups, where our proposed model is dominating over others. TT2VFin+p and TT2VFin+M have no change in performance, but TT2VFin+p+M cannot maintain its previous performance. In general, our proposed model uses 2 to 3 times fewer parameters than others but has roughly 14.23% greater performance (by averaging all the metrics differences between base and RNN and LSTM).
@@ V Conclusion
In this paper, we investigated the application of deep learning for stock price prediction, which is a challenging time series forecasting problem in the financial industry. Notably, we represent a novel methodology for combining correlation features, aiming to better capture inter-stock relationships and temporal dynamics. We studied different markets by selecting correlation features to enhance the accuracy in predicting multiple stock prices and developed a novel neural network architecture that combines the Encoder of the Transformer with the Time2Vec model. Through careful testing and analysis, we have found that our model performs exceptionally well, better than standard models such as LSTM and RNN, and encoding methods like positional encoding from Transformer. In most cases, our multi-feature proposed model is better than a single-feature model, that verifies the importance and potential of feature selection and aggregation. In conclusion, our study demonstrated that techniques like Time2Vec and Encoder are promising future approaches combined with multi-source feature selection. We also believe that our approach can be successfully adopted and generalized to related financial and other time series prediction problems, and can achieve even better results for more refined data sources (e.g., having the intra-day historical data).